# FinTagging: An LLM-ready Benchmark for Extracting and Structuring Financial Information

### v1

@@ 3.1 Task Formulation
$f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$ (1)
@@
Inspired by information extraction and alignment works wu2019scalable ; wadden2019entity ; mukherjee2022ectsum ; wang2022conditional and after discussed with financial reporting specialists, we formulated FinTagging into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type. Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
Financial Numeric Identification (FinNI).
[…]
@@
$f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$ (2)
@@
[…]
$f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$ (3)
@@
3.2 Raw Data Collection
We collected 30 annual 10-K reports filed in 2024 by publicly listed companies from SEC  [^6] https://www.sec.gov/, as summarized in Table 2. There are a total of 81,325 facts, of which 69,451 are linked to standard taxonomy tags and 11,874 to SEC extension tags. Furthermore, using the BeautifulSoup tool to parse these reports, we identified 76,835 narrative sentences (approximately 16 million characters) and 5,450 financial tables. The companies’ information is shown in Appendix C. These reports follow the XBRL standard. We include all sections that may contain XBRL-tagged content to ensure comprehensive coverage.
Table 2: The statistic of collected financial reports.
Report type Period #Company #Sentence #Char #Table #Std Tag #Custom Tag #Total Tag
Information 10-k 2024-02-13 to 2025-02-13 30 76,835 16,595,283 5,450 69,451 11,874 81,325
3.3 FinTagging Data Annotation
To construct the FinTagging benchmark, we annotated a total of 3,354 sentences and 3,245 table sequences extracted from the narrative and tabular content of collected financial reports, as summarized in Table 3. Among them, 1,627 textual instances and 2,014 table sequences contain XBRL tagging information (positive), while the remaining 1,727 textual and 1,231 table instances do not (negative). On average, textual inputs contain 82.55 tokens, whereas table sequences are significantly longer, averaging 999.74 tokens. Based on this, we further construct two subtask-specific datasets, FinNI-eval and FinCL-eval, with the following steps. The statistics for these two datasets are shown in Table 4.
@@ 3.3 FinTagging Data Annotation
Table 3: The statistical information for the original dataset in our benchmark. Using the “cl100k_base” tokenizer to calculate tokens ($\pm$ standard deviation).
Sentence Positive 1,627 $82.55\pm 66.39$ $1.27\pm 1.84$ $1.27\pm 1.84$ 52,740 2,288
Negative 1,727
Table Positive 2,014 $999.74\pm 993.37$ $14.89\pm 23.50$ $14.89\pm 23.50$
Negative 1,231
For narrative sentences, we retain only textual segments that exceed 20 characters in length and contain numerical terms. For table cell sequences, we use specific labels such as <table>, <tr>, <td>, and <th> to preserve layout and structural information. Tables that contain only headers without any numerical values are considered invalid and thus excluded from our benchmark.
3.3.1 Curation of Numeric Entities and US-GAAP Concepts
Building on the initial filtering, we further refined the sentence and table set to retain only high-quality samples with gold-standard annotations of numerical entity types and their associated US-GAAP concepts. As shown in Table 3, we obtained 52,740 annotated numerical entities spanning 5 entity types, linked to a total of 2,288 unique US-GAAP concepts.
Specifically, we predefine the five of the most prevalent item types, including monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType. We first excluded sentences and tables lacking US-GAAP labels. Then we conducted a statistical analysis of the entity types contained in the data, as summarized in Figure 2, to identify the most frequent types.
Figure 2: The statistic of numerical entity type.
From Figure 2, we identified 11 numerical entity types across the financial reports of 30 companies. The top five types, based on frequency and coverage, appear in all reports. Sentences or tables containing these types are labeled as positive instances; others are treated as negative.
@@ 3.3.1 Curation of Numeric Entities and US-GAAP Concepts
Table 4: Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Using the “cl100k_base” tokenizer to calculate tokens ($\pm$ standard deviation).
Dataset #Instance Avg Input Tokens Max Input Tokens Avg Output Tokens Max Output Tokens
FinNI-eval 6,599 $986.11\pm 835.52$ 13,205 $138.28\pm 302.92$ 5,465
FinCL-eval 52,572 $61.11\pm 44.60$ 750 $25.00\pm 213.77$ 37
3.3.2 FinNI-eval Dataset Construction
@@
Based on the annotation procedure described in Section 3.3.1, we construct the FinNI-eval dataset, comprising 6,599 samples derived from the annotated sentences and tables. Specifically, the dataset comprises two components: the input block and the answer block. The input block includes a task instruction and the input content. The instruction defines the FinNI task by specifying the responsibilities of the LLMs, the definitions of entity types, and the rules to be followed. The answer block is formatted as a list aligned with the input block. The answer is a JSON list of all identified entity values and types or an empty JSON list. The specific prompt template is shown in the Appendix H.1.
@@
3.3.3 FinCL-eval Dataset Construction
Following Section 3.1, we further construct a FinCL-eval dataset that includes 52,572 query-answer pairs for numerical entity normalization. In addition, we build a US-GAAP taxonomy database containing 17,388 unique financial concepts. To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3.1. Each query consists of a numerical entity, its corresponding entity type, and its surrounding context, while the answer is the associated US-GAAP concept.
@@ 3.3.3 FinCL-eval Dataset Construction
3.4 Evaluation
3.4.1 Evaluation Framework
[…]
Figure 3: A unified evaluation framework on FinTagging benchmark.
FinNI Evaluation
@@ 3.4.1 Evaluation Framework
We use pair-level metrics to evaluate the extraction performance on FinNI. The objective is to evaluate the LLM’s ability to distinguish between structured and unstructured contexts and to extract financial values based on its semantic understanding and domain-specific knowledge.
FinCL Evaluation
[…]
$f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$ (4)
3.4.2 Evaluation Models
@@
Our goal is to evaluate the foundational capabilities of SOTA LLMs on the FinTagging benchmark, to better understand their strengths and limitations in financial tagging. To this end, we evaluate models across three categories using our proposed evaluation framework, as detailed in Appendix E. These include: (1) General-purpose closed-source LLM: GPT-4o model hurst2024gpt ; (2) General-purpose open-source LLMs including DeepSeek-V3 liu2024deepseek , DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability , Qwen2.5-0.5B-Instruct qwen2 , Qwen2.5-1.5B-Instruct qwen2 , Qwen2.5-14B-Instruct qwen2 , Llama-3.2-3B-Instruct grattafiori2024llama , Llama-3.1-8B-Instruct grattafiori2024llama , and gemma-2-27b-it gemma_2024 models; and (3) Domain-specific financial LLM: Fino1-8B qian2025fino1 model. In addition, we compare these LLMs with strong PLMs, including BERT-large DBLP:journals/corr/abs-1810-04805 , FinBERT araci2019finbert , and SECBERT loukas2022finer .
3.4.3 Evaluation setting

### v2

@@ 3.1. Task Formulation
(1) $f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$
@@
Inspired by information extraction and alignment works (Wu et al., 2019; Wadden et al., 2019; Mukherjee et al., 2022; Wang et al., 2022) and after discussed with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type. Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
Financial Numeric Identification (FinNI)
The first subtask of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: focuses on identifying numerical values in a financial document and assigning each a coarse-grained value data type. This corresponds to detecting the Fact and Type components of each triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$ defined in the overall task.
@@
(2) $f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$
@@
Building on the output of the FinNI subtask, the goal of FinCL is to semantically ground each identified numerical entity $e$ by linking it to a concept $\hat{c}$ in a predefined financial taxonomy, which is the Tag component of each triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$ defined in the overall task. Formally, we define the mapping as:
(3) $f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$
@@
3.2. Raw Data Collection
We compiled 142 annual 10-K reports filed in 2023 and 2024 by publicly listed companies from the SEC [^6] https://www.sec.gov/, as shown in Table 2. To ensure diversity, the collection covers all 11 major industry sectors, with about 95% of companies distributed across 31 U.S. states and the remaining 5% based outside the United States. Using BeautifulSoup, we parsed these reports and extracted 319,893 narrative sentences (roughly 76 million characters) along with 21,576 financial tables. Company-level details are provided in Appendix D.2. Each report follows the SEC’s XBRL filing standard, meaning that the underlying instance, schema, and linkbase documents already contain machine-readable tags (ground-truth tags) that link textual and tabular values to US-GAAP taxonomy concepts. We included all sections that may reference or embed XBRL-tagged content to ensure comprehensive coverage. For the justification of our benchmark granularity, please see the Appendix E.
Table 2. Statistics of collected financial reports.
Item Value
Report type 10-K
Period 2023-02-13 to 2025-02-13
#Company 142
#Covered sector 11
#Covered jurisdiction 31 States + outside US
#Sentence 319,893
#Char 75,748,949
#Table 21,576
3.3. Data Curation
To construct the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark and ensure both efficiency and fairness in evaluation, we designed a two-stage filtering and annotation pipeline. This pipeline systematically processes the raw 10-K filings to retain representative, high-quality instances that collectively cover all 142 companies, five valid entity types, and every numerical US-GAAP concept observed in the corpus. The overall workflow is illustrated in Algorithms 1 and 2.
Stage 1: Filtering Raw Instances (Algorithm 1)
The first stage aims to remove noisy or redundant samples while ensuring comprehensive coverage of both entity types and US-GAAP concepts. Each narrative sentence or table instance is automatically examined by a rule-based parser that checks if it contains numeric entities tagged with valid US-GAAP concepts. The parser performs schema-level consistency checks based on the taxonomy definitions to ensure tagging correctness. Instances without valid entities are classified as NEG (negative). Those containing valid entities are treated as positive and further divided into two categories: (1) POS_SELECT, which introduces at least one new US-GAAP concept not yet covered; and (2) POS_CANDIDATE, which contains only previously covered concepts. This design ensures that each concept appears at least once in the dataset while minimizing redundancy across companies and reports.
Figure 2. Statistics of numerical entity types.
To determine which entity types should be considered valid, we first profiled all XBRL numeric item types that appeared in the raw corpus and summarized the results in Figure 2. For each type, we calculated both its cross-company coverage (the number of companies in which it occurs) and its overall frequency (the proportion of all annotated numeric entities belonging to that type). The figure shows a clear long-tailed distribution: a total of eleven numeric item types were identified, but only a few dominate in both coverage and frequency. The five types that consistently appear across nearly all companies and collectively account for the majority of entities are monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType. These five types form the label space for all subsequent stages. Under this definition, Algorithm 1 treats an instance as a potential positive only if it contains at least one US-GAAP–tagged entity belonging to this label space; otherwise it is assigned to NEG.
To maintain structural fidelity, additional validity rules are applied. For narrative texts, we retain only segments longer than 20 characters that contain numerical tokens. For tables, we preserve the <table>, <tr>, <td>, and <th> tags to ensure layout integrity, while tables with headers only (i.e., without numeric values) are excluded.
Input: 10-K reports; US-GAAP taxonomy; 5 valid entity types
Output: POS_SELECT, POS_CANDIDATE, NEG
Initialize: an empty set of covered concepts; empty lists for POS_SELECT, POS_CANDIDATE, and NEG;
foreach instance (texts or tables) in 10-K reports do
  Check if the instance contains numeric entities that are both: (i) tagged with a US-GAAP concept, and (ii) of a valid entity type;
 if no such entities are found then
     Add the instance to NEG;
 else
     Identify all US-GAAP concepts in this instance;
     if the instance introduces at least one new concept not yet covered then
        Add the instance to POS_SELECT and update the covered concepts set;
    else
        Add the instance to POS_CANDIDATE;
return {POS_SELECT, POS_CANDIDATE, NEG};
Algorithm 1 Screening of Text and Table Instances
Stage 2: Company-Capped Sampling and Minimal Cover (Algorithm 2)
The second stage refines the positive pool by enforcing both per-company balance and concept coverage completeness. For each modality (text or table), we first perform company-capped seeding: from each company’s POS_SELECT, up to $K=10$ positive instances are retained as the seed set, while the remaining positives are moved into a candidate pool. We then mark all US-GAAP concepts covered by the seed and identify those still uncovered. Finally, a greedy minimal-cover procedure is executed: candidates are iteratively selected according to the number of uncovered concepts they can cover, with ties broken by shorter text length, fewer entities, or earlier order. Each selected candidate is added to the final set $P^{*}$, and the process continues until all concepts are covered or the candidate pool is exhausted. The resulting positive set $P^{*}$ thus maintains full US-GAAP concept coverage, balanced company representation, and reduced redundancy, forming a compact yet coverage-complete dataset.
Input: POS_SELECT grouped by company (run separately for text and table) from algorithm 1;
   All concepts observed in the POS_SELECT;
   Cap $K=10$ (maximum positives per company for seeding)
Output: Final positive instances $P^{*}$ and leftover instances
Seed selection. For each company, keep up to $K$ positives as the seed set. Put the rest into a new candidate pool.
Check coverage. Mark all concepts already covered by the seed. Identify which concepts are still uncovered.
Greedy covering. While uncovered concepts remain:
  1. (1)
    For each candidate, count how many uncovered concepts it can cover.
  2. (2)
    Pick the candidate with the highest count (ties broken by shorter text, fewer entities, or earlier order).
  3. (3)
    Add the candidate to the final positives, update the covered concepts, and remove it from the pool.
Finalize. The final positives $P^{*}$ are the union of the seed and the greedy additions. The leftover candidates form the candidate set.
Algorithm 2 Company-Capped Seeding + Minimal Cover (for Text and Table)
Final Dataset and Validation
After two-stage filtering, we obtained 3,084 textual sentences and 3,067 table sequences from financial filings (Table 3). Among them, 2,054 text and 2,045 table instances are positives containing valid XBRL tagging information, while 1,030 text and 1,022 table instances are negatives. Textual inputs contain on average $96.25$ tokens, whereas table sequences are much longer ($1,204.72$ tokens on average). Overall, the dataset includes 62,668 annotated entities linked to 3,953 unique US-GAAP concepts across the five valid entity types.
Based on this foundation, we further derive two subtask-specific datasets: FinNI-eval, targeting numerical entity identification, and FinCL-eval, targeting concept linking. Both are the first datasets in the financial domain to emphasize structured information extraction and concept-level alignment. Their detailed statistics are reported in Table 4.
To verify annotation reliability, two domain experts independently reviewed 100 randomly sampled entries. The results indicate a 96% raw agreement and a Cohen’s Kappa of 81.13, confirming the robustness and reproducibility of our semi-automated annotation pipeline. The validation guideline is provided in Appendix F, and a complete unfiltered version of the benchmark is included in Appendix D.1.
@@ 3.3. Data Curation
Table 3. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
Item Sentence Table
Positive instances 2,054 2,045
Negative instances 1,030 1,022
Avg. Tokens/S $96.25\pm 72.75$ $1,204.72\pm 1,162.41$
Avg. Entities/S $1.88\pm 2.11$ $18.54\pm 29.14$
Avg. Concepts/S $1.88\pm 2.11$ $18.54\pm 29.14$
Total Entities 62,668
Entity Types 5
Unique Concepts 3,953
@@
Table 4. Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
Item FinNI-eval FinCL-eval
#Instance 6,151 62,668
Avg. Input Tokens $1101.50\pm 992.00$ $60.43\pm 48.25$
Max Input Tokens 18,831 750
Avg. Output Tokens $175.24\pm 375.90$ $12.82\pm 5.09$
Max Output Tokens 7,813 41
3.4. FinNI-eval Dataset Construction
@@
Based on the annotation procedure described in Section 3.3, we construct the FinNI-eval dataset, comprising 6,151 instances derived from the annotated sentences and tables. Specifically, the dataset comprises two components: the input block and the answer block. The input block includes a task instruction and the input content. The instruction defines the FinNI task by specifying the responsibilities of the LLMs, the definitions of entity types, and the rules to be followed. The answer block is formatted as a list aligned with the input block. The answer is a JSON list of all identified entity values and types or an empty JSON list. The specific prompt template is shown in Figure 9 in the Appendix.
@@
3.5. FinCL-eval Dataset Construction
Following the formulation of the FinCL subtask in Section 3.1, we further construct a FinCL-eval dataset that includes 62,668 query-answer pairs for numerical entity normalization. In addition, we built a US-GAAP taxonomy database containing 17,688 unique financial concepts. To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3. Each query consists of a numerical entity, its corresponding entity type, and its surrounding context, while the answer is the associated US-GAAP concept.
@@ 3.5. FinCL-eval Dataset Construction
3.6. Evaluation
3.6.1. Evaluation Framework
We propose a unified evaluation framework for the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark to assess LLM performance. As shown in Figure 3, given a financial report $D$, the framework reformulates tagging into two subtasks, FinNI and FinCL, to generate structured triplets $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$. This design jointly evaluates an LLM’s zero-shot ability in fact extraction and concept alignment. For fair comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 (Sokolova and Lapalme, 2009). FinNI is evaluated using pair-level metrics, while FinCL uses accuracy. Full metric definitions are provided in Appendix G.
Figure 3. A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. Note: The Sentence is drawn from financial report $D$, and the Query provides an example that incorporates the entity 4.9. The black arrows show the overall framework flow, green arrows denote the FinNI process, red arrows indicate the FinCL process, red bidirectional arrows represent retrieval interactions with the taxonomy, and red dashed arrows mark the recomposed input between the query and retrieved candidates.
@@ 3.6.1. Evaluation Framework
FinNI Evaluation: We use pair-level metrics to evaluate the extraction performance on FinNI. The objective is to evaluate the LLM’s ability to distinguish between structured and unstructured contexts and to extract financial values based on its semantic understanding and domain-specific knowledge.
FinCL Evaluation: To assess whether LLMs can perform fine-grained XBRL entity normalization, we reformulate the task as a retrieval-reranking problem, where LLMs are used to disambiguate and select the most appropriate taxonomy concept (reranking stage) from a reduced candidate set (retrieval stage), avoiding the impracticality of direct multi-thousand-way classification (prompt template used for reranking in the FinCL subtask is shown in Figure 10 in Appendix).
To this end, we first generate embeddings for each taxonomy concept using text-embedding-3-small [^7] https://platform.openai.com/docs/models/text-embedding-3-small, and retrieve the top-$h$ candidate tags from the US-GAAP taxonomy $\mathcal{T}$ based on semantic similarity. Therefore, the LLMs are required to rerank the retrieved candidates by leveraging deeper contextual understanding to select the best-matched tag $\hat{c}\in\mathcal{T}_{h}$, formalized as:
(4) $f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$
3.6.2. Evaluation Models
@@
Our goal is to evaluate the foundational capabilities of SOTA LLMs on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, to better understand their strengths and limitations in financial tagging. To this end, we evaluate models across three categories using our proposed evaluation framework, as detailed in Table 13. These include: (1) one general-purpose closed-source LLM: GPT-4o model (Hurst et al., 2024); (2) eleven general-purpose open-source LLMs including DeepSeek-V3 (Liu et al., 2024), DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025), Qwen3-32B (Team, 2025), Qwen3-14B (Team, 2025), Qwen3-1.7B (Team, 2025), Qwen3-0.6B (Team, 2025), Llama-4-Scout-17B-16E-Instruct (Meta, 2025), Llama-3.3-70B-Instruct (Grattafiori et al., 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Llama-3.2-3B-Instruct (Grattafiori et al., 2024), and gemma-2-27b-it (Team, 2024) models; and (3) one domain-specific financial LLM: Fino1-8B (Qian et al., 2025) model. In addition, we compare these LLMs with strong PLMs, including BERT-large (Devlin et al., 2018), FinBERT (Araci, 2019), and SECBERT (Loukas et al., 2022).
3.6.3. Evaluation setting